Mikaela1509:

what is 3/2 times 36/45 as a fraction my brother really needs help

LKylie:

I didnt get the question

LKylie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mikaela1509 like what is it as a mixed fraction? \(\color{#0cbb34}{\text{End of Quote}}\) 36/45 as a mixed fraction?

AZ:

Just multiply the numerators and denominators \(\dfrac{a}{b}\times \dfrac{c}{d} = \dfrac{a\times b}{c \times d}\) So for your question \(\dfrac{3}{2} \times \dfrac{36}{45} = \dfrac{3\times 36}{2\times 45} = ??\) what is 3 * 36 and what is 2 * 45 use a calculator